OpenStudy (anonymous):

The functions f and g are defined by f(x) = 2x + 1 and g(x) = 3 for all numbers x. What is the least value of c for which f(c) = g(c)?

OpenStudy (jack1):

when f(x) = g(x) f(x) = g(c) 2c + 1 = 3??? yeah...?

OpenStudy (jack1):

so 2c = 3-1 = 2... so c = 1 for all values of f(x) = g(x) ... ? i think...?

OpenStudy (anonymous):

yeahh i did this same..but answer is -2

OpenStudy (anonymous):

its asking for the least value...

OpenStudy (anonymous):

but i dont know how -2 is ans

OpenStudy (jack1):

hmmm... im still lost... how could c = -2 -2 x 2 +1 = -3...?

OpenStudy (anonymous):

hehehe..yeah thts what m too confused about :-)

OpenStudy (jack1):

\[\large -3 \neq 3\] :-/ nuh, sorry mann, i got nothin hay

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

please can anybody help me out..to get this question right?

OpenStudy (anas.p):

jack is right... can u make sure the question is correct.

OpenStudy (anonymous):

yeahh i checked it thrice!!!

OpenStudy (anonymous):

www.number2.com this the link ..i got this ques from here..(GRE quant)

OpenStudy (anas.p):

sorry man.. i still think jack is right...